Please explain how the simplifying works to get from:
D/3 = 4 + D/9,
TO:
3D = 36 + D
Also, where could I find more about this type of problem?

Answer by usyim88hk(157)

Just multiply the original equation by 9
(D/3)(9) = 4(9)+(D/9)(9)
3D = 36+D